# Bohao Wu   Qingyun Wang   Yue Guo

### v1

@@ Appendix A Setup
Dataset
We use data from  Guo et al. (2024), which includes four personalized classification tasks: term familiarity and the need for additional information in the form of a definition, background, or example. Each is framed as a binary classification problem. The dataset contains 1,100 annotations per task, collected from 11 researchers who each labeled 100 out-of-domain abstracts, selecting 10 terms per abstract. The dataset also includes researcher metadata, publication records, and related background information; further details can be found in the original paper. We follow the original 60/20/20 split for training, validation, and testing to ensure direct comparability with their reported baselines.
@@
To evaluate the performance of our personalized jargon identification models, we focus on the task of predicting binary familiarity labels (0 for familiar, 1 for unfamiliar) for entities extracted from research paper abstracts. Our primary evaluation metric is the F-1 score. However, During our initial baseline model selection phase, we observed that some models struggled to consistently produce the required binary label lists without additional text or nonsensical information. To account for this, we introduced the Effective F-1 Score. This metric incorporates the "Mismatch Rate" – the proportion of model outputs that did not conform to the expected binary label format. The Effective F-1 Score is calculated as follows:
[…]
@@
To establish a robust foundation for our personalized jargon identification task using Parameter Efficient Fine-Tuning (PEFT), we first selected a suitable open-source Large Language Model (LLM) as our baseline.
We evaluated several state-of-the-art options, including Llama-2 (Touvron et al., 2023), Llama-3 (Dubey et al., 2024), Mistral v0.3 (Jiang et al., 2023), and Qwen-2.5 (Yang et al., 2024; Team, 2024), considering both their base and instruction-tuned versions in 4-bit and full precision.
The results of this initial evaluation (depicted in Figure 6) revealed significant performance variations in terms of both F-1 score and mismatch rate. While Qwen-2.5 7B Instruct 4bit achieved the highest Effective F-1 Score (0.54), and Llama-3.1 8B 4bit exhibited the lowest mismatch rate (0.5%), we ultimately selected Llama-3.1 8B Instruct 4bit as our baseline for subsequent fine-tuning experiments. This decision was based on its robust performance (Effective F-1 Score of 0.49) and its demonstrated ability to follow instructions with minimal mismatches, suggesting a strong potential for effective adaptation through PEFT for our personalized jargon identification task.
Appendix B Analysis
Does the familiarity model generalize over other personalized tasks?
@@ Appendix B Analysis
Model Improvement Analysis in Terms of Individual Annotators
@@
Taking annotator #4 as the object, a qualitative analysis of the missed and falsely detected jargon reveals several interesting patterns. The baseline model (trained with 1% training set), while showing some capability in jargon detection, struggled with terms that exhibit a combination of characteristics. Firstly, it frequently failed to identify the terms as jargons that are relatively short and composed of common words but carry highly specific meanings within a particular domain. Examples include ’Radial curves’ (Materials Science), ’Op-amp’ (Physics), and ’Domains’ (Geology). These terms, due to their brevity and seemingly ordinary components, may have been harder for the baseline to differentiate from general language use. Secondly, the baseline model had difficulty with multi-word terms where the meaning is not a straightforward combination of the individual words, but rather a more nuanced concept. This is evident in its failure to identify ’Bayesian optimal mechanism’ (Economics), ’Riemannian framework’ (Materials Science), ’Bose-Einstein condensate’ (Physics), ’Psychometric properties’ (Economics), ’Dialectical quality’ (Philosophy), ’Explanatory account’ (Linguistics), ’Long-range ordered coupling’ (Physics, Materials Science), and ’Qualitative spatio-temporal inferences’ (Psychology). In these cases, the model may have lacked the ability to capture the semantic relationships and contextual dependencies necessary for accurate identification. Thirdly, the baseline also missed acronyms like ’CW-SSIM’ (Agricultural And Food Sciences), ’MANOVA’ (Education), and ’ARMAX model’ (Business, Engineering). Acronyms often present a challenge due to their condensed nature and lack of explicit semantic clues. Finally, there were instances where the jargon term spans multiple disciplines, such as ’Monolayers’ (Engineering, Biology), ’Peri-implant bone density’ (Materials Science, Medicine, Biology), and ’Regulatory mechanisms’ (Biology, Environmental Science), which might have added to the difficulty. While the improved model demonstrated a higher F1 score, indicative of better overall performance, it exhibited a tendency to produce more false positives. These false positives included terms like ’Savitzky-Golay (SG) filter’ (Environmental Science), ’Meta-analyses’ (Medicine), ’Post-test’ (Education), ’Quantitative research’ (Education), and ’Content analysis’ (Medicine). This suggests that the improved model, in its attempt to capture a broader range of jargon, may be more sensitive to terms that share some characteristics with jargon but are more commonly used or understood. This could indicate a trade-off where the improved model sacrifices some precision for increased recall, potentially overgeneralizing in certain contexts. Specifically, the improved model seems to be more prone to misclassifying statistical or methodological terms (e.g., ’Meta-analyses’, ’Post-test’, ’Quantitative research’) as jargon, possibly due to their frequent appearance in academic contexts, even when they are relatively well-understood within the broader research community.
Model Analysis in Terms of Jargon Domain
@@
In this study, the best two models are selected, which are 100% TS with vanilla and nearest abstract (NAb) prompting. When comparing the two models across the ’Art’ and ’Philosophy’ domains, a nuanced performance profile emerges. In the ’Art’ domain, vanilla model exhibits a higher false positive rate, incorrectly identifying terms like ’Reactions’ and ’Stylistic’ as jargon, which NAb model correctly classifies. This suggests that vanilla model may be overly sensitive to terms that, while potentially used in art-related contexts, also have broader, common usage.
Conversely, in the ’Philosophy’ domain, NAb model faces challenges in both precision and recall. It exhibits a higher false positive rate, misclassifying terms such as ’Structural constraints’, ’Analytic philosophers’, and ’Argument Facets’. This indicates a tendency to over-identify common philosophical terms as highly specialized jargon. Furthermore, NAb model also demonstrates lower recall in the Philosophy domain, failing to detect several jargon terms, including ’Computational argumentation’, ’Corpus with 320 arguments’, ’Nonmonotonic inference methods’, ’Super-knotty rope’, ’Super-knot’, and ’Dialectical quality’. These terms represent complex philosophical concepts that NAb model struggles to recognize as domain-specific jargon.
Models F-1 Score $\uparrow$ Recall Precision
Oracle Classifier 71.9±1.7 76.0±2.1 68.2±2.1
GPT-4 64.2±1.5 98.7±0.5 47.6±1.6
1%TS 53.5±2.8 56.0±4.5 51.5±3.2
10%TS 63.6±2.9 59.8±5.1 69.0±1.7
Unsuper 54.6±5.1 77.5±7.6 45.0±0.8
Unsuper+10%TS 77.0±1.1 78.9±2.6 75.4±0.7
100%TS 77.9±1.2 77.8±2.2 78.0±0.5
100%TS+MD 76.8±1.1 76.1±2.9 77.7±1.0
100%TS+Profile 76.6±1.0 73.7±2.0 79.9±1.6
100%TS+NAn 72.1±3.0 70.3±4.1 74.5±2.7
100%TS+NAb 77.8±1.1 78.3±1.7 77.5±2.5
Table 1: Evaluation of the finetuned models on the test dataset. If not specified, the finetuning and inference are performed with vanilla prompting. Acronym: ‘TS’-traing set, ‘MD’-metadata, ‘NAn’-nearest annotator, ‘NAb’-nearest abstract.

### v2

@@ Appendix A Setup
To evaluate the performance of our personalized jargon identification models, we focus on predicting binary familiarity labels (0 for familiar, 1 for unfamiliar) for entities extracted from research paper abstracts. Our primary evaluation metric is the F-1 score. However, during our initial baseline model selection phase, we observed that some models struggled to consistently produce the required binary label lists without additional text or nonsensical information. To account for this, we introduced the Effective F-1 Score. This metric incorporates the “Mismatch Rate”, the proportion of model outputs that did not conform to the expected binary label format. The Effective F-1 Score is calculated as follows:
$\text{eff. F-1 score}=(1-\text{Mis. rate})\times\text{F-1 score.}$
@@
To establish a robust foundation for our personalized jargon identification task using Parameter Efficient Fine-Tuning (PEFT), we first selected a suitable open-source Large Language Model (LLM) as our baseline. We evaluated several state-of-the-art options, including Llama-2 (Touvron et al., 2023), Llama-3 (Dubey et al., 2024), Mistral v0.3 (Jiang et al., 2023), and Qwen-2.5 (Yang et al., 2024; Team, 2024), considering both their base and instruction-tuned versions in 4-bit and full precision.
For our implementation, we use the unsloth library with all parameters set to their default values, including is_bfloat16_ supported. The results of this initial evaluation (depicted in Figure 5) revealed significant performance variations in terms of both F-1 score and mismatch rate. While Qwen-2.5 7B Instruct 4bit achieved the highest Effective F-1 Score (0.54), and Llama-3.1 8B 4bit exhibited the lowest mismatch rate (0.5%), we ultimately selected Llama-3.1 8B Instruct 4bit as our baseline for subsequent fine-tuning experiments. This decision was based on its robust performance (Effective F-1 Score of 0.49) and its demonstrated ability to follow instructions with minimal mismatches, suggesting a strong potential for effective adaptation through PEFT for our personalized jargon identification task.
Appendix B Additional Analysis
B.1 Does the familiarity model generalize over other personalized tasks?
@@
B.2 Model Improvement Analysis in Terms of Individual Annotators
@@
Taking annotator #4 as the object, a qualitative analysis of the missed and falsely detected jargon reveals several interesting patterns. The baseline model (trained with 1% training set), while showing some capability in jargon detection, struggled with terms that exhibit a combination of characteristics. Firstly, it frequently failed to identify the terms as jargon that are relatively short and composed of common words but carry highly specific meanings within a particular domain. Examples include ‘Radial curves’ (Materials Science), ‘Op-amp’ (Physics), and ‘Domains’ (Geology). These terms, due to their brevity and seemingly ordinary components, may have been harder for the baseline to differentiate from general language use. Secondly, the baseline model had difficulty with multi-word terms where the meaning is not a straightforward combination of the individual words, but rather a more nuanced concept. This is evident in its failure to identify ‘Bayesian optimal mechanism’ (Economics), ‘Riemannian framework’ (Materials Science), ‘Bose-Einstein condensate’ (Physics), ‘Psychometric properties’ (Economics), ‘Dialectical quality’ (Philosophy), ‘Explanatory account’ (Linguistics), ‘Long-range ordered coupling’ (Physics, Materials Science), and ‘Qualitative spatio-temporal inferences’ (Psychology). In these cases, the model may have lacked the ability to capture the semantic relationships and contextual dependencies necessary for accurate identification. Thirdly, the baseline also missed acronyms like ‘CW-SSIM’ (Agricultural And Food Sciences), ‘MANOVA’ (Education), and ‘ARMAX model’ (Business, Engineering). Acronyms often present a challenge due to their condensed nature and lack of explicit semantic clues. Finally, there were instances where the jargon term spans multiple disciplines, such as ‘Monolayers’ (Engineering, Biology), ‘Peri-implant bone density’ (Materials Science, Medicine, Biology), and ‘Regulatory mechanisms’ (Biology, Environmental Science), which might have added to the difficulty. While the improved model demonstrated a higher F1 score, indicative of better overall performance, it exhibited a tendency to produce more false positives. These false positives included terms like ‘Savitzky-Golay (SG) filter’ (Environmental Science), ‘Meta-analyses’ (Medicine), ‘Post-test’ (Education), ‘Quantitative research’ (Education), and ‘Content analysis’ (Medicine). This suggests that the improved model, in its attempt to capture a broader range of jargon, may be more sensitive to terms that share some characteristics with jargon but are more commonly used or understood. This could indicate a trade-off where the improved model sacrifices some precision for increased recall, potentially overgeneralizing in certain contexts. Specifically, the improved model appears to be more prone to misclassifying statistical or methodological terms (e.g., ‘Meta-analyses’, ’Post-test’, ‘Quantitative research’) as jargon, possibly due to their frequent occurrence in academic contexts, even when they are relatively well understood within the broader research community.
B.3 Model Analysis in Terms of Jargon Domain
@@
In this study, the two best models are selected, which are 100% TS with vanilla and nearest abstract (NAb) prompting. When comparing the two models across the ‘Art’ and ‘Philosophy’ domains, a nuanced performance profile emerges. In the ‘Art’ domain, the vanilla model exhibits a higher false positive rate, incorrectly identifying terms like ‘Reactions’ and ‘Stylistic’ as jargon, whereas the NAb model correctly classifies them. This suggests that vanilla model may be overly sensitive to terms that, while potentially used in art-related contexts, also have broader, common usage.
Conversely, in the ’Philosophy’ domain, NAb model faces challenges in both precision and recall. It exhibits a higher false positive rate, misclassifying terms such as ’Structural constraints’, ’Analytic philosophers’, and ’Argument Facets’. This indicates a tendency to over-identify common philosophical terms as highly specialized jargon. Furthermore, NAb model also demonstrates lower recall in the Philosophy domain, failing to detect several jargon terms, including ‘Computational argumentation’, ‘Corpus with 320 arguments’, ‘Nonmonotonic inference methods’, ‘Super-knotty rope’, ‘Super-knot’, and ‘Dialectical quality’. These terms represent complex philosophical concepts that the NAb model struggles to recognize as domain-specific jargon.